QUESTION 2 White eyes is an X-linked recessive mutation and curly wings is an autosomal dominant mutation in Drosophila. What proportion of red eyed, straight winged males is expected in the F2 starting with a true breeding white eyed, non-curly female mating with a true breeding male mutant only for curly. b3/16 Ос.О el e-1/32 QUESTION 3 Red-Green colorblindness is a X-linked recessive disorder. Huntingtons is an autosomal dominant disorder. Susans father is color blind and has Huntingtons disease and is heterozygous for the Huntingtons allele. Her mother is normal and has no history of colorblindness or Huntingtons in her family. Her brother has developed Huntingtons but is not colorblind. Her husband Greg is healthy and has no incidence of either disorder in his family. Susan is not colorblind and has not yet be tested to see if she has inherited the Huntingtons allele. What is the chance that a child of Susan and Greg will be a girl with normal vision but having Huntingtons? O a. 1/4 b. 1/64 1/8 el e1/32
Answer #1

2. Answer: Option D is correct.
White eyes = Recessive = Xw
Red eyes = Dominant = XW
Straight wings = Recessive = c
Curly wings = Dominant = C

Parental cross: XwXw cc X XWY CC
F1 progeny: XWXwCc XwYCc
F1 selfing: XWXwCc X XwYCc
Required progeny: XWYcc
Frequency = 1/4 X 1/4
= 1/16

3. Answer: Option C is correct.
Colorblindenss = Recessive = Xw
Normal vision = Dominant = XW
Normal allele = Recessive = h
Huntington's allele = Dominant = H

Susan's father = XwYHh
Susan's mother = XWXWhh
Susan's brother = XWYHh
Susan's husband = XWYhh
Susan = XWXwHh or XWXwhh

Parental cross: XwXw cc X XWYhh
Probability for a girl = 1/2
Probability for a girl with a normal vision = 1
Probability for a girl with Huntington allele = 1/4

Combined frequency = 1/2 X 1 X 1/4
= 1/8